A man starts with an initial velocity of 3.50 m/s and accelerates for a distance of 205 m over 28.7s. What is the acceleration of the man?


1
Expert's answer
2022-05-16T16:06:51-0400

"s=vt+\frac{at^2}2,"

"a=\frac{2s}{t^2}-\frac{2 v}t=0.25~\frac m{s^2}."